OpenStudy (anonymous):

Which of the following ordered pairs is a solution to the inequality y <= -2x + 10? (0, -4) (1, 13) (6, -1) (3, 8)

OpenStudy (anonymous):

@Kanwar245

OpenStudy (anonymous):

(0,-4)

OpenStudy (anonymous):

so A or C?

OpenStudy (anonymous):

Which of the following is not an equivalent form of the compound inequality x + 18 <= 11 and x + 18 >= 3 A) A number line with a closed circle on -15, shading to the left, and a closed circle on -7, shading to the right. B) 11 >= x + 18 >= 3 C) -7 >= x >= -15 D) A number line with a closed circle on -15, a closed circle on -7, and shading in between.